OpenStudy (18jonea):

The conservation of mechanical energy says that the initial potential energy of an object plus the final potential energy of the object equals the initial kinetic energy of the object plus the final kinetic energy of the object. true false

OpenStudy (sooobored):

As i mentioned in the previous post the conservation of energy says that the total energy (kinetic and potential) of 1 state must be equal to the total energy of the 2nd state

OpenStudy (sooobored):

so KE1+PE1 = constant = KE2+PE2

OpenStudy (sooobored):

what your problem has written is incorrect PE1+PE2= KE1+KE2

OpenStudy (18jonea):

ok false

OpenStudy (18jonea):

@sooobored

OpenStudy (sooobored):

yes...

OpenStudy (18jonea):

thank you!